# Enhancing Precision of Signal Correction in PVES Experiments: The Impact of Bayesian Analysis on the Results of the QWeak and MOLLER Experiments

### v1

@@ V Conclusion
This study highlights the advantages of Bayesian analysis in enhancing the precision of signal correction by improving the estimation of asymmetry uncertainties in PVES experiments, as evidenced by the reanalysis of the QWeak inelastic data and the methodology proposed for the MOLLER experiment. Our results indicate that Bayesian methods provide more consistent uncertainty estimates than traditional Monte Carlo approaches.
Moreover, transitioning to a Bayesian approach offers a robust framework that not only refines the accuracy of signal corrections but also deepens our understanding of the underlying physical processes. The promising outcomes from the QWeak and MOLLER experiments establish a foundation for applying Bayesian analysis to upcoming high-precision parity-violation experiments such as P2 and SoLID. This method is anticipated to be equally effective in addressing the unique challenges presented by these experiments.

### v2

@@ V Conclusion
In this study, the reanalysis of the QWeak inelastic data using the Bayesian method, compared to the previously employed Many-Worlds Monte Carlo Minimization approach, provided insights into the precision and reliability of both techniques. The Bayesian analysis produced smaller uncertainties and more stable estimations, offering a closer fit to the measured asymmetry values. In contrast, the Monte Carlo method resulted in larger uncertainties and greater variations across detectors, likely due to the incorrect assumption of Gaussian distributions for non-Gaussian inputs and the approach to quantify the uncertainties. Despite these differences, both methods yielded central values consistent within their respective uncertainties, ensuring that the physics conclusions of the original study remained unchanged. The analysis of the MOLLER experiment using Bayesian inference demonstrated an agreement between the simulated input asymmetry values and the extracted output values within their respective uncertainty ranges. The comparative analysis of fitted versus mock asymmetry values further validated the Bayesian approach, with the residuals centered around zero and exhibiting small RMS values. Furthermore, the correlation analysis in both experiments revealed insights into the interdependencies of asymmetry components, with QWeak showing expected correlations between longitudinal and transverse asymmetries, while the MOLLER experiment’s setup effectively separated detector signals, resulting in no correlations between asymmetry parameters. The promising outcomes from the QWeak and MOLLER experiments establish a foundation for applying Bayesian analysis to upcoming high-precision parity-violation experiments such as P2 and SoLID. This method is anticipated to be equally effective in addressing the unique challenges presented by these experiments.